# Quench dynamics via recursion method and dynamical quantum phase transitions

### v2

@@ Abstract
The recursion method provides a powerful framework for studying quantum many-body dynamics in the Lanczos basis recursively constructed within the Krylov space of operators. Recently, it has been demonstrated that the recursion method, when supplemented by the universal operator growth hypothesis, can effectively compute autocorrelation functions and transport coefficients at infinite temperature. We extend the scope of the recursion method to far-from-equilibrium quench dynamics. We apply the method to spin systems in one, two, and three spatial dimensions. In one dimension we find an evidence that the convergence of the method is limited in time by the dynamical quantum phase transition.
@@
Formally, the output of the recursion method is the Heisenberg operator of an observable. This suggests that the scope of the method can be much broader than infinite-temperature correlation functions and should cover, in particular, the out-of-equilibrium dynamics after quantum quenches. Addressing the later in systems of dimension higher than one is known to be highly challenging, with limited body of results obtained to date Schmitt and Heyl (2020); Li et al. (2020); Dziarmaga (2022); Bernier and Agarwal (2025); Begušić and Chan (2025); Loizeau et al. (2025a, b). Given the success of the recursion method in computing two-dimensional autocorrelation function, it seems tempting to apply it to quench dynamics.
Here we undertake exactly this attempt. The paper is organized as follows. We start from briefly introducing the recursion method and specifying it for addressing quench dynamics. Then we apply it to a paradigmatic one-dimensional mixed-field Ising model. The results are benchmarked by essentially exact numerical simulations. We demonstrate that the recursion method has a convergence domain limited in time by the dynamical quantum phase transition. Then results for the transverse-field Ising model on the square and cubic lattices are presented. We conclude by discussing implications of our results and open questions raised.
[…]
@@
$b_{n}$ $\simeq\frac{\alpha\,n}{\log n}+\gamma+\frac{(-1)^{n}\alpha}{\log n+c}$ ${\rm for}~{}~{}1{\rm D},$ (4)
$b_{n}$ $\simeq\alpha\,n+\gamma+\frac{(-1)^{n}\alpha}{\log n+c}$ ${\rm for}~{}~{}2,3{\rm D}.$ (5)
@@
where $\langle O^{n}\rangle\equiv\operatorname{\mbox{tr}}\rho_{0}O^{n}$. We compute $O^{n}$ and $b_{n}$ for $n$ up to $n_{\rm max}$ symbolically directly in the thermodynamic limit, as described in Uskov and Lychkovskiy (2024). $\langle O^{n}\rangle$ is also computed symbolically. As for functions $\phi^{n}_{t}$, they are obtained by numerically solving coupled linear differential equations (3) with the tridiagonal matrix ${\mathcal{L}}$ truncated at some large $k$ (we use $k=1000$ in the examples below). The entries of this matrix for $n>n_{\rm max}$ are obtained from extrapolation of the $n_{\rm max}$ known Lanczos coefficients according to extrapolation formulae (4), (5).
@@
Figure 1: Time evolution of the (a) Loschmidt rate and (b)–(d) total polarization along $z$ direction in the 1D Ising model (7) with $h_{x}=h_{z}=1$. The initial state is a product translation-invariant state (9) with the polarization $(p_{x},p_{y},p_{z})$ indicated on each plot. Orange line – essentially exact results of the numerical time evolution (see text for details). Blue line - results of the recursion method. The Loschmidt rate features multiple cusp singularities manifesting dynamical quantum phase transitions. The results of the recursion method depart from the exact numerical results at time $t_{*}$ close to the time $t_{c}$ of the second cusp (red dashed vertical line).
Figure 2: Scaling of the estimated breakdown time $t_{*}$ with the number $n_{\rm max}$ of explicitly know Lanczos coefficients for various initial polarizations. $t_{*}$ is defined as the time when the recursion method result departs from the essentially exact numerical result by 0.005. The dashed horizontal line marks the time $t_{c}$ of the dynamical quantum phase transition, see Fig. 1 (a). Clearly, $t_{*}$ exhibits a tendency to saturate around $t_{c}$.
@@
One can see from Fig. 1 that the recursion method excellently reproduces the actual $\langle A_{t}\rangle$ initially, but then abruptly breaks down. We observe that this breakdown is associated with a dynamical quantum phase transition (DQPT) Pollmann et al. (2010); Heyl et al. (2013); Karrasch and Schuricht (2013); Heyl (2018) – a phenomenon by which the Loschmidt rate function $\lambda(t)=N^{-1}\,\log\big{|}\langle\Psi_{0}|e^{-i\,t\,H}|\Psi_{0}\rangle\big{|}^{2}$ features singularities in the thermodynamic limit.111 There is an ambiguity in the meaning of the term “dynamical (quantum) phase transition.” Historically, it was introduced in a quite different context, referring to a transition between distinct phases of matter that unfolds over time due to variations in the Hamiltonian Schützhold et al. (2006); Eckstein et al. (2009); Uhlmann et al. (2010); Garrahan and Lesanovsky (2010); Diehl et al. (2010); Sciolla and Biroli (2010) – a definition more closely aligned with the conventional understanding of phase transitions. In contrast, the definition we adopt was introduced later Heyl et al. (2013) and is based on a formal analogy between the Loschmidt rate and the free energy. The breakdown time $t_{*}$ is very close to the time $t_{c}$ of one of the singularities (specifically, the second singularity), see Fig. 1.
More precisely, the breakdown time $t_{*}$ grows with the number $n_{\rm max}$ of available explicit Lanczos coefficient, but this growth saturates at $t_{c}$, as shown in Fig. 2. This strengthens our belief that the observed association between the breakdown of the recursion method and the DQPT is not accidental.
Remarkably, while the recursion method breaks down at a finite time $t_{*}$, at large times it seems to typically converge to the actual equilibrium value of the observable, as illustrated in Fig. 3 (a). Note that the one-dimensional transverse-field Ising model is known for its unusually slow thermalization Bañuls et al. (2011); Lin and Motrunich (2017); Peng and Cui (2022); Pathak (2024) which is seen in Fig. 3 (a).
@@
We compute $23$ and $12$ nested commutators, respectively. Time evolution of the magnetization is presented in Fig. 3 (b),(c). The results of the recursion method are compared to the outcomes of the sparse Pauli dynamics method Begušić and Chan (2025); Beg .
One can see that the agreement is good initially but abruptly worsens at some threshold time. We are not able to decisively resolve the origin of this discrepancy at the moment. Of course, the DQPT is again the main suspect. DQPT is known to occur in the 2D Ising model, although with a softer logarithmic singularity Heyl (2018). However, we are not able to compute the critical time for this DQPT for two- and three-dimensional models.
@@
In one dimension, benchmarking is straightforward due to the availability of essentially exact numerical simulations. By comparing the recursion method to these simulations, we have observed that it yields accurate results up to a certain threshold time, beyond which it abruptly breaks down.
The breakdown indicates that the expansion (6) of a Heisenberg observable in the Lanczos basis has a finite convergence domain, at least for one-dimensional systems. This fact may limit the scope of concepts and techniques based on this expansion Parker et al. (2019); Chu et al. (2024); Loizeau et al. (2025c).
Notably, the breakdown time turns out to closely coincide with the critical time of a dynamical quantum phase transition (DQPT), suggesting that DQPT fundamentally limits the applicability of the recursion method, at least in one dimension. The underlying mechanism linking DQPT to the breakdown of the recursion method remains entirely unknown and represents a pressing open question. The potential subtlety of this mechanism is highlighted by our observation that the breakdown time coincides not with the first, but with the second (in time) dynamical critical point, implying that not every DQPT necessarily disrupts the recursion method.
In contrast to the one-dimensional case, benchmark data in two and three dimensions is scarce. We have compared our results to those obtained from the sparse Pauli dynamics method Begušić and Chan (2025); Beg and found that the two approaches diverge abruptly after a certain threshold time, similar to what is observed in one dimension. However, due to the lack of available data on DQPTs in the higher-dimensional models we considered, we cannot unambiguously attribute this divergence to DQPTs. Determining whether DQPTs fundamentally constrain the recursion method’s convergence in higher dimensions remains an important open question.
Finally, in one dimension, we have observed that at long times, the recursion method’s outcome coincides with the time-averaged value of the observable, which, in thermalizing systems, is equal to the equilibrium value. If this observation proves to be correct and generic, it would imply an unusual scenario: the recursion method remains valid at both short and long times but fails at intermediate times. Investigating this phenomenon further presents another promising direction for future research.
Note added.  While this manuscript has been in preparation, a closely related study has appeared Loizeau et al. (2025c). The latter paper also addresses quench dynamics in one-dimensional spin systems by means of the recursion method. It does not report the finite convergence domain of the method or any association with the DQPT.
@@ Acknowledgements.
Acknowledgments.  We sincerely thank Alexey Rubtsov for pointing to the dynamical phase transition as a potential cause of the recursion method breakdown. We are grateful to Tomislav Begušić and Garnet Kin-Lic Chan for discussing their paper Begušić and Chan (2025) and our manuscript, and particularly for providing unpublished data Beg on the 2D Ising model. We thank Berislav Buča and Nicolas Loizeau for useful remarks and discussions. Finally, numerous discussions with Alexander Teretenkov and Nikolay Il’in are gratefully acknowledged. This work was supported by the Russian Science Foundation under grant № 24-22-00331, https://rscf.ru/en/project/24-22-00331/

### v3

@@ Abstract
The recursion method provides a powerful framework for studying quantum many-body dynamics in the Lanczos basis recursively constructed within the Krylov space of operators. Recently, it has been demonstrated that the recursion method, when supplemented by the universal operator growth hypothesis, can effectively compute autocorrelation functions and transport coefficients at infinite temperature. We extend the scope of the recursion method to far-from-equilibrium quench dynamics. We apply the method to spin systems in one, two, and three spatial dimensions. In one dimension its usefulness is limited: although it remains accurate at moderate times, it eventually experiences an abrupt breakdown and subsequently yields results that deviate strongly from the true dynamics. In contrast, in two and three dimensions the method proves far more effective, providing a reasonably accurate description of the evolution across all time scales – from the initial transient regime to thermalization.
@@
Formally, the output of the recursion method is the Heisenberg operator of an observable. This suggests that the scope of the method can be much broader than infinite-temperature correlation functions and should cover, in particular, the out-of-equilibrium dynamics after quantum quenches. Addressing the later in systems of dimension higher than one is known to be highly challenging, with limited body of results obtained to date Schmitt and Heyl (2020); Li et al. (2020); Dziarmaga (2022); Bernier and Agarwal (2025); Sinibaldi et al. (2024); Gravina et al. (2025); Pavešić et al. (2025); Chen et al. (2025); Begušić and Chan (2025); Loizeau et al. (2025a, b); Rudolph et al. (2025). Given the success of the recursion method in computing two-dimensional autocorrelation function, it seems tempting to apply it to quench dynamics.
Here we undertake exactly this attempt. The paper is organized as follows. We start from briefly introducing the recursion method and specifying it for addressing quench dynamics. Then we apply it to a paradigmatic one-dimensional mixed-field Ising model. The results are benchmarked by essentially exact numerical simulations. Then results for the transverse-field Ising model on the square and cubic lattices are presented. We conclude by discussing implications of our results and open questions raised.
Recursion method for quench dynamics.  Consider an observable given by a self-adjoint Shrödinger operator $A$. The same observable in the Heisenberg representation reads $A_{t}=e^{itH}A\,e^{-itH}$, where $H$ is the Hamiltonian of the system. The Heisenberg operator $A_{t}$ evolves within the Krylov space spanned by operators $A,{\mathcal{L}}A,{\mathcal{L}}^{2}A,\dots$, where ${\mathcal{L}}\equiv[H,\bullet]$ is the superoperator referred to as Liouvillian. $A_{t}$ satisfies the Heisenberg equation of motion $\partial_{t}A_{t}=i\,{\mathcal{L}}A_{t}$. Within the recursion method Viswanath and Müller (2008), this equation is addressed in the orthonormal Lanczos basis $\{O^{0},O^{1},O^{2},\dots\}$ constructed as follows: $O^{0}=A/\|A\|$, $A^{1}={\mathcal{L}}O^{0}$, $b_{1}=\|A^{1}\|$, $O^{1}=b_{1}^{-1}\,A^{1}$ and $A^{n}={\cal L}\,O^{n-1}-\,b_{n-1}O^{n-2}$, $b_{n}=\|A^{n}\|$, $O^{n}=b_{n}^{-1}A^{n}$ for $n\geq 2$. Here a scalar product in the space of operators is introduced according to $\big(A|B\big)\equiv{\cal N}\,\operatorname{\mbox{tr}}\big(A^{\dagger}B\big)$, where $\cal N$} is a normalization constant. For systems of $N$ spins $1/2$ and translation-invariant observables considered here it is convenient to choose ${\cal N}=N^{-1}\,2^{-N}$. The scalar product entails the norm $\|A\|=\sqrt{(A|A)}$. The coefficients $b_{n}$ are referred to as Lanczos coefficients. In the Lanczos basis the Liouvillian has a tridiagonal form,
@@
$b_{n}$ $\simeq\frac{\alpha\,n}{\log n}+\gamma+\frac{(-1)^{n}\tilde{\alpha}}{\log n+c}$ ${\rm for}~~1{\rm D},$ (4)
$b_{n}$ $\simeq\alpha\,n+\gamma+\frac{(-1)^{n}\tilde{\alpha}}{\log n+c}$ ${\rm for}~~2,3{\rm D}.$ (5)
@@
where $\langle O^{n}\rangle\equiv\operatorname{\mbox{tr}}\rho_{0}O^{n}$. We compute $O^{n}$ and $b_{n}$ for $n$ up to $n_{\rm max}$ symbolically directly in the thermodynamic limit, as described in Uskov and Lychkovskiy (2024). $\langle O^{n}\rangle$ is also computed symbolically. Final numerical values of $b_{n}$ are obtained with the use of rational arithmetic, which avoids the numerical instability of the Lanczos algorithm Eckseler et al. (2025). As for functions $\phi^{n}_{t}$, they are obtained by numerically solving coupled linear differential equations (3) with the tridiagonal matrix $\mathcal{L}$} truncated at some large $k$ (we use $k=1000$ in the examples below). The entries of this matrix for $n>n_{\rm max}$ are obtained from extrapolation of the $n_{\rm max}$ known Lanczos coefficients according to extrapolation formulae (4), (5).
@@
Figure 1: Time evolution of the total polarization along $z$ direction in the 1D Ising model (7) with $h_{x}=h_{z}=1$. The initial state is a product translation-invariant state (9) with the polarization $(p_{x},p_{y},p_{z})$ indicated on each plot. Orange line – essentially exact results of the numerical time evolution (see text for details). Blue line - results of the recursion method. The recursion method reproduces initial evolution but abruptly departs from the exact numerical result at the breakdown time $t_{*}$ (vertical dashed line). This time appears to be independent of the initial state.
@@
One can see from Fig. 1 that the recursion method excellently reproduces the actual $\langle A_{t}\rangle$ initially, but then abruptly breaks down at the breakdown time $t=t_{*}$. After this breakdown, the recursion method result strongly deviates from the actual value of the observable. We note that the breakdown time appears to be independent of the initial state.
We attribute this breakdown of the recursion method to the slow thermalization typical of one-dimensional systems. In particular, the mixed-field Ising model is well known for its anomalous thermalization behavior Bañuls et al. (2011); Lin and Motrunich (2017); Kormos et al. (2017); Peng and Cui (2022); Pathak (2024).
In the previous version of this manuscript Shirokov et al. (2025), we linked the breakdown of the recursion method to a dynamical quantum phase transition (DQPT), motivated by an apparent coincidence between the breakdown time and the DQPT time. A more careful analysis shows that this coincidence is not universal and does not hold across most of the parameter space. We therefore retract our earlier claim that a DQPT is the cause of, or a universal precursor to, the breakdown of the recursion method in one dimension. Additional data and discussion on this point are provided in the Supplement.
Remarkably, at large times the recursion method result typically converges to the actual equilibrium value of the observable, as illustrated in Fig. 2 (a). This convergence may indicate that the terms discarded by the truncation in Eq. (2), while significantly affecting the transient dynamics, do not contribute much in the limit of $t\rightarrow\infty$.
@@
We compute $23$ and $17$ nested commutators, respectively. Time evolution of the magnetization is presented in Fig. 2 (b),(c). The results of the recursion method are compared to the outcomes of the sparse Pauli dynamics (SPD) method Begušić and Chan (2025); Beg . Up to the maximal time accessible to the SPD, the agreement with our results is excellent at early times and remains reasonably good thereafter. In contrast to the SPD, however, the recursion method yields predictions all the way to equilibration.
@@
In one dimension, benchmarking is straightforward due to the availability of essentially exact numerical simulations. By comparing the recursion method to these simulations, we have observed that it yields accurate results up to a certain threshold time, beyond which it abruptly breaks down. Despite this breakdown, the recursion method result eventually equilibrates close to the actual equilibrium value.
The breakdown indicates the slow convergence of the expansion (6) of a Heisenberg observable in the Lanczos basis. This slow convergence may limit the scope of concepts and techniques Parker et al. (2019); Chu et al. (2024); Loizeau et al. (2025c); Alishahiha and Vasli (2025); Rabinovici et al. (2025) built on this expansion in one dimension.
In contrast to the one-dimensional case, benchmark data in two – and especially three – dimensions remain scarce. We have compared our results with those obtained using the SPD method Begušić and Chan (2025); Beg and find agreement ranging from excellent to reasonable up to the maximal times accessible to SPD. Importantly, the recursion method is essentially unrestricted in time and continues to yield reasonable results all the way to equilibration.
Note added.  While this manuscript has been in preparation, a closely related study has appeared Loizeau et al. (2025c). The latter paper addresses quench dynamics in one-dimensional spin systems by means of the recursion method.
@@ Acknowledgements.
Acknowledgments.  We are grateful to Tomislav Begušić and Garnet Kin-Lic Chan for discussing their paper Begušić and Chan (2025) and our manuscript, and particularly for providing unpublished data Beg on the 2D Ising model. We thank Berislav Buča and Nicolas Loizeau for useful remarks and discussions. Finally, numerous discussions with Alexander Teretenkov and Nikolay Il’in are gratefully acknowledged. This work was supported by the Russian Science Foundation under grant № 24-22-00331, https://rscf.ru/en/project/24-22-00331/
@@ Appendix A Supplementary material
Note that in the previous version of the manuscript Shirokov et al. (2025), for the 3D Ising model, we provided 12 moments and Lanczos coefficients for the non-translation-invariant observable $\sigma^{z}_{j}$. Here we provide 17 moments and Lanczos coefficients for the respective translation-invariant observable. While the values are different in the two cases, the output of the recursion method is consistent whenever the initial state is translation-invariant.
@@
Appendix E S4. Recursion method breakdown and dynamical quantum phase transition in one dimension
Figure S3: Time evolution of the total polarization along $z$ direction in the 1D Ising model (7) for various values of $h_{z}$ and $h_{x}$. The initial state is a product translation-invariant state (9) polarized in the $z$-direction. Orange line – essentially exact results of the numerical time evolution (see text for details). Blue line – results of the recursion method. Pink line – the rate function $\lambda(t)$ whose cusps mark the dynamical quantum phase transitions.
In the previous version of this manuscript Shirokov et al. (2025), we have attributed the breakdown of the recursion method to a dynamical quantum phase transition (DQPT) Pollmann et al. (2010); Heyl et al. (2013); Karrasch and Schuricht (2013); Heyl (2018) – a phenomenon by which the Loschmidt rate function $\lambda(t)=-N^{-1}\,\log\big|\langle\Psi_{0}|e^{-i\,t\,H}|\Psi_{0}\rangle\big|^{2}$ features singularities in the thermodynamic limit. [^1] There is an ambiguity in the meaning of the term “dynamical (quantum) phase transition.” Historically, it was introduced in a quite different context, referring to a transition between distinct phases of matter that unfolds over time due to variations in the Hamiltonian Schützhold et al. (2006); Eckstein et al. (2009); Uhlmann et al. (2010); Garrahan and Lesanovsky (2010); Diehl et al. (2010); Sciolla and Biroli (2010) – a definition more closely aligned with the conventional understanding of phase transitions. In contrast, the definition we adopt was introduced later Heyl et al. (2013) and is based on a formal analogy between the Loschmidt rate and the free energy.
Here, we have examined this conjecture more thoroughly across a range of model parameters. The results, shown in Fig. S3, indicate that, in general, $t_{*}$ and $t_{c}$ do not coincide. The apparent agreement occurs, within uncertainties, only for $h_{x}\simeq h_{z}$. We therefore retract our earlier strong claim regarding a relation between the recursion method breakdown and the DQPT, leaving further investigation of this connection to future work. For the sake of this future work, we point to a paper Takahashi (2025) on DQPT and Lanczos method in the Schrödinger representation.